OpenStudy (anonymous):

Can someone help me on a last question!? Please Ill give a medal. "Prove that a line parallel to one side of a triangle divides the other two sides proportionally. Be sure to create and name the appropriate geometric figures. "

ganeshie8 (ganeshie8):

draw it out

OpenStudy (anonymous):

I dont know how and the picture didn't came with the question.

ganeshie8 (ganeshie8):

its okay, lets draw a rough picture :) just a triangle and a line parallel to one of its sides

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

u got some clue ? or still need some assistance ha ? :)

OpenStudy (anonymous):

I feel if you do the shape like this i would understand it more. because i know the geometric figure 

ganeshie8 (ganeshie8):

that works too :) lets stick to ur diagram

OpenStudy (anonymous):

Ok great :)

ganeshie8 (ganeshie8):

1. \(\angle D \cong \angle A\) 1. Corresponding angles 2. \(\angle B \cong \angle B\) 2. By reflexive property 3. \(\triangle ABC \sim \triangle DBE\) 3. By AA similarity

ganeshie8 (ganeshie8):

half part of the proof is done. see if you can digest above...

ganeshie8 (ganeshie8):

we have just proved that both triangles are similar; next we will take ratios and conclude the proof

OpenStudy (anonymous):

AD/AB = BE/BC

OpenStudy (anonymous):

I have a feeling I did this wrong?

ganeshie8 (ganeshie8):

yes slightly wrong. u need to honor the sides-belonging-to-respective-triangles

ganeshie8 (ganeshie8):

wat that means is simply this :- sides belonging to same triangle must be on TOP or BOTTOM.

ganeshie8 (ganeshie8):

AD/AB = BE/BC this is wrong cuz, AD and BE are on top, but they dont belong to same triangle!

OpenStudy (anonymous):

So AB and BC are suppose to be on top.

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

you got it ! write the correct proportion now :)

OpenStudy (anonymous):

AB/AD = BC/BE

OpenStudy (anonymous):

Like this?

ganeshie8 (ganeshie8):